# Stationary solutions to stochastic 3D Euler equations in Hölder space

### v1

@@ 4.6.2. Stationary phase Lemma
Hence, (3.8d), (3.8e) and (3.8f) hold ture at the level $q+1$.
@@ 4.7. Inductive estimates for the energy
To conclude the proof of Proposition 3.2, we shall verify (3.9) holds at level $q+1$. To this end, we define
$\delta E(t):=\Big{|}e(t)(1-\delta_{q+2})-\mathbf{E}\|v_{q+1}(t)+{z_{q+1}}(t)\|_{L^{2}}^{2}\Big{|}.$
@@
$\delta E(t)\leqslant$ $\frac{1}{4}\delta_{q+2}e(t).$ (4.52)
@@
By the definition of $\zeta_{q}$ in Section 4.3, we find
@@
which shall be estimated. Let us begin with the bound of the first term on the right hand side of (4.53). Similar as the computation of oscillation error (4.29) in Section 4.5, by (4.13) we have for any $t\in[k,k+1]$, $k\in\mathbb{Z}$
$w_{q+1}^{(p)}\otimes w_{q+1}^{(p)}$
@@
Taking the trace on both sides of the equation and using the fact that $\mathring{R}_{\ell}$ is traceless, we deduce for $t\in[k,k+1]$, $k\in\mathbb{Z}$
@@
In the following, we estimate each term separately, by the choice of parameters, we find
@@
which requires $b^{2}\beta<1$ and $a$ sufficiently large to absorb the constant. Using (3.8f) on $\mathring{R}_{q}$, $\operatorname{supp}\varphi_{\ell}\subset[0,\ell]$ and $c_{R}<c_{*}$ we deduce for $t\in{\mathbb{R}}$
@@
For the third term in (4.54), we use (3.3), (3.7), (3.8b) and (3.8c) to obtain
@@
where we used $14b^{2}\beta+2\beta<$ 1 to have $\lambda_{q}^{-\frac{1}{2}+\beta}<\lambda_{q}^{-2b^{2}\beta}$ and $\lambda_{q}^{-\frac{5}{6}+\beta}<\lambda_{q}^{-2b^{2}\beta}$, and we $a$ was chosen sufficiently large to absorb the constant. For the last term in (4.54), recalling the definition of $\eta_{k,j}$ and $a_{(\xi)}$ in (4.12), we know that for any fixed $t\in[k,k+1]$, the sum with respect to $j,j^{\prime}$ is actually the finite sum of at most two terms. Hence, applying (4.40) with $n=m$ and using the estimates (4.14), (4.15) and (C.7) we obtain
@@
where we used $b^{3}\gamma+2b^{2}\beta<1$ and $5m+5<bm$ to have $\lambda_{q}^{4m+2+(m+2)b^{3}\gamma}<\lambda_{q}^{bm-2b^{2}\beta}$, and we chose $a$ sufficiently large to absorb the constant. This completes the bound for (4.54). Going back to (4.53), it remains to control
@@
where we used $b\geqslant 7$, $14b^{3}\gamma+4b^{2}\beta<$ 1 to have $\lambda_{q}^{2b^{3}\gamma-\frac{1}{2}}<\lambda_{q}^{-2b^{2}\beta}$, $\lambda_{q}^{4+3b^{3}\gamma}<\lambda_{q}^{b-2b^{2}\beta}$ and $\lambda_{q}^{8+4b^{3}\gamma}<\lambda_{q}^{2b-2b^{2}\beta}$, and $a$ was chosen sufficiently large to absorb the constant. Next, we use (3.3), (3.7), (3.8b), (4.6) and (4.33) to obtain
@@
where we used $2b\beta+\beta<\frac{\gamma\sigma}{8}$ to have $\lambda_{q}^{b\beta+2b^{2}\beta}<\lambda_{q}^{\frac{\gamma\sigma}{8}b}$, and used $14b^{2}\beta+2b\beta<$ 1, $3b\beta<\gamma$ to have $\lambda_{q}^{b\beta-\frac{1}{2}}<\lambda_{q}^{-2b^{2}\beta}$ and $\lambda_{q}^{-b^{2}\gamma+b\beta+b\gamma}<\lambda_{q}^{-2b^{2}\beta}$, and $a$ was chosen large enough to absorb the constant. For the last term of (4.55), we recall (cf.(4.19)) that $w_{q+1}$ may be written as the curl of a vector field as
@@
where we used $b\geqslant 7$ and $12b^{2}\beta+b^{3}\gamma<$ 1 to have $\lambda_{q}^{b^{3}\gamma+\frac{3}{2}}<\lambda_{q}^{b-2b^{2}\beta}$ and $\lambda_{q}^{b^{3}\gamma+b\gamma}<\lambda_{q}^{b-2b^{2}\beta}$, and $a$ was chosen sufficiently large to absorb the constant. Summing over the above estimates, we obtain
@@
Hence, the proof of Proposition 3.2 is completed. ∎
@@ 5. Stationary solutions to the stochastic Euler equations
We recall that the trajectory space defined in Section 1.1 is $\mathcal{T}=C({\mathbb{R}};C^{\kappa})\times C({\mathbb{R}};C^{\kappa})$ and we take $\kappa=\frac{\vartheta}{2}$ in this section. The corresponding shifts $S_{t}$, $t\in{\mathbb{R}}$, on trajectories are given by
@@
The notion of stationary solution was introduced in Definition 1.4. Our main result of this section is existence of stationary solutions as limits of ergodic averages of solutions constructed in Section 3. This in particular implies their non-uniqueness.
@@
$\sup_{s\in{\mathbb{R}}}\mathbf{E}\|u(\cdot+s)\|^{2r}_{C^{\beta^{\prime}}([-N,N],C^{\beta^{\prime\prime}}_{x})}\lesssim N.$
Indeed, reviewing the proof of Theorem 3.3, and by interpolation we deduce for some $\beta^{\prime}\in(0,\frac{\vartheta}{2})$ and $\beta^{\prime\prime}\in(\frac{\vartheta}{2},\vartheta)$ satisfying $\beta^{\prime}+\beta^{\prime\prime}<\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot b^{5}}},\frac{1}{21{\cdot}b^{4}}\}}$
@@
where we used $(\frac{1}{4}-\beta^{\prime})(1-\beta^{{}^{\prime\prime}})>\frac{1}{16}$, $b\geqslant 7$ and $\beta^{\prime\prime}<{\mathord{{\rm min}}}{\{\frac{\sigma}{4b},1\}}$ in the last two inequalities. Hence we obtain $\lim_{q\rightarrow\infty}z_{q}=z$ in $L^{p}(\Omega,C^{\beta^{\prime}}(\mathbb{R};C^{\beta^{\prime\prime}}))$ for any $p\geqslant 1$. Then letting $u=v+z$, we obtain an $(\mathcal{F}_{t})_{t\in\mathbb{R}}$-adapted analytically weak solution $u$ to (1.1) and
@@
Hence we obtain
$\sup_{s\in{\mathbb{R}}}\mathbf{E}\|u(\cdot+s)\|^{2r}_{C^{\beta^{\prime}}([-N,N],C^{\beta^{\prime\prime}}_{x})}$ $\leqslant\sup_{s\in{\mathbb{R}}}\sum_{i=-N}^{N-1}\mathbf{E}\|u(\cdot+s)\|^{2r}_{C^{\beta^{\prime}}([i,i+1],C^{\beta^{\prime\prime}}_{x})}$
$\leqslant 2N\sup_{t\in\mathbb{R}}\mathbf{E}\|u\|^{2r}_{C^{\beta^{\prime}}([t,t+1],C^{\beta^{\prime\prime}}_{x})}\lesssim N.$
Now, we define the ergodic averages of the solution $(u,B)$ as the probability measures on the space of trajectories ${\mathcal{T}}$
@@
For $R_{N}>0$, $N\in\mathbb{N}$, by Arzelà–Ascoli Theorem and the above assertion, we note that the set
@@
is relatively compact in $C({\mathbb{R}};C^{\frac{\vartheta}{2}})$ since $\frac{\vartheta}{2}<\beta^{\prime\prime}$. As a consequence, we deduce that the time shifts $S_{t}u$, $t\in{\mathbb{R}}$, are tight on $C({\mathbb{R}};C^{\frac{\vartheta}{2}})$. Since $S_{t}B$ is a Wiener process for every $t\in{\mathbb{R}}$, the law of $S_{t}B$ does not change with $t\in{\mathbb{R}}$ and is tight. Accordingly, for any $\varepsilon>0$ there is a compact set $\bar{K}_{\varepsilon}$ in ${\mathcal{T}}$ such that
$\sup_{t\in{\mathbb{R}}}{\mathbf{P}}(S_{t}(u,B)\in\bar{K}^{c}_{\varepsilon})<\varepsilon.$
@@
$\nu_{T}(\bar{K}_{\varepsilon}^{c})=$ $\frac{1}{T}\int_{0}^{T}\mathbf{P}(S_{t}(u,B)\in\bar{K}^{c}_{\varepsilon}){\mathord{{\rm d}}}t<\varepsilon$
and therefore there is a weakly converging subsequence of the probability measures $\nu_{T}$, $T\geqslant 0$. That is, there is a subsequence $T_{n}\to\infty$ and $\nu\in{\mathcal{P}}({\mathcal{T}})$ such that $\nu_{T_{n}}\to\nu$ weakly in ${\mathcal{P}}({\mathcal{T}})$. Define the set
$A$ =$=\Big{\{}(u,B)\in{\mathcal{T}};\ \langle u(t),\psi\rangle+\int_{s}^{t}\langle\mathord{{\rm div}}(u\otimes u),\psi\rangle{\mathord{{\rm d}}}r$
$\qquad\qquad=\langle u(s),\psi\rangle+\langle B(t)-B(s),\psi\rangle,\quad\forall\psi\in C^{\infty}(\mathbb{T}^{3};{\mathbb{R}}^{3}),\ \mathord{{\rm div}}\psi=0,\ t\geqslant s\Big{\}}.$
Since $(u,B)$ in the statement of the theorem satisfies the equation, we have for all $t\in{\mathbb{R}}$
${\mathcal{L}}[S_{t}(u,B)](A)=1.$
Hence, also $\nu_{T_{n}}(A)=1$ for all $n\in\mathbb{N}$. By Jakubowski–Skorokhod representation theorem, there is a probability space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbf{P}})$ with a sequence of random variables $(\tilde{u}^{n},\tilde{B}^{n})$, $n\in\mathbb{N}$, such that ${\mathcal{L}}[\tilde{u}^{n},\tilde{B}^{n}]={\nu_{T_{n}}}$ and $(\tilde{u}^{n},\tilde{B}^{n})$ satisfy equation (1.1) on ${\mathbb{R}}$. By (3.12) we know
$\tilde{\mathbf{E}}\|\tilde{u}^{n}(t)\|_{L^{2}}^{2}=\frac{1}{T_{n}}\int_{0}^{T_{n}}\mathbf{E}\|{S_{s}u(t)}\|_{L^{2}}^{2}{\mathord{{\rm d}}}s=K.$ (5.2)
By (3.11)
$\sup_{n}\tilde{\mathbf{E}}\|\tilde{u}^{n}(t)\|_{L^{2}}^{2r}=\sup_{n}\frac{1}{T_{n}}\int_{0}^{T_{n}}\mathbf{E}\|S_{s}u(t)\|_{L^{2}}^{2r}{\mathord{{\rm d}}}s<\infty.$ (5.3)
Moreover, there is a random variable $(\tilde{u},\tilde{B})$ having the law ${\mathcal{L}}[\tilde{u},\tilde{B}]=\nu$ so that
$(\tilde{u}^{n},\tilde{B}^{n})\to(\tilde{u},\tilde{B})\qquad\tilde{\mathbf{P}}\text{-a.s. in }\mathcal{T}.$
Thus, we can pass to the limit in the equation to deduce that $\nu$ is a law of a solution on ${\mathbb{R}}$. The shift invariance follows from the same argument as in [BFH20e, Lemma 5.2]. Namely, it holds for $G\in C_{b}({\mathcal{T}})$ and $r\in{\mathbb{R}}$
$\int_{{\mathcal{T}}}G\circ S_{r}(u,B){\mathord{{\rm d}}}\nu(u,B)=\int_{{\mathcal{T}}}G(u,B){\mathord{{\rm d}}}\nu(u,B).$
Finally, (5.1) follows from (5.2) and (5.3). ∎
Using the above result and choosing different $K$, the second claim in Theorem 1.5 follows.
@@ Appendix 1 A Beltrami waves
In this part we recall the Beltrami waves from [BV19, Section 5.4] whcih is adpated to the convex integration scheme in Proposition 3.2. We point out that the construction is entirely deterministic, that is, none of the functions below depends on $\omega$. Let us begin with the definition of Beltrami waves. Given $\xi\in\mathbb{S}^{2}\cap\mathbb{Q}^{3}$, let $A_{\xi}\in\mathbb{S}^{2}\cap\mathbb{Q}^{3}$ obey
@@
is $\mathbb{T}^{3}$ periodic, divergence free, and is an eigenfunction of the $\mathrm{curl}$ operator with eigenvalue $\lambda$. That is, $W_{(\xi)}$ is a complex Beltrami plane wave. The following lemma states a useful property for linear combinations of complex Beltrami plane waves, which can be found in [BV19, Proposition 5.5].
@@
The key point of the construction is that the abundance of Beltrami waves allows to find several such flows $v$ with the property that
@@
equals a prescribed symmetric matrix $R$. Indeed we will need to select these flows so as to depend smoothly on the matrix $R$, at least when $R$ belongs to a neighborhood of the identity matrix. In view of (A.3), such selection is made possible by the following lemma.
@@
$\Lambda_{\alpha}\subset\mathbb{S}^{2}\cap\mathbb{Q}^{3}\qquad\alpha\in\{0,1\},$
@@
$\gamma_{\xi}^{(\alpha)}\in C^{\infty}(B_{c_{*}}(\mathrm{Id}))\qquad\alpha\in\{0,1\},\xi\in\Lambda_{\alpha},$

### v3

@@ 5.2.7. Estimate on the commutator error
Hence, (3.8d), (3.8e) and (3.8f) hold at the level $q+1$.
@@ 5.3. Inductive estimates for the energy
To conclude the proof of Proposition 3.3, we shall verify (3.9) holds at the level $q+1$.
@@
$\Big{|}e(t)(1-\delta_{q+2})-\mathbf{E}\|v_{q+1}(t)+{z_{q+1}}(t)\|_{L^{2}}^{2}\Big{|}\leqslant$ $\frac{1}{4}\delta_{q+2}e(t).$ (5.37)
@@
By the definition of $\zeta_{q}$ in (4.7), we have
@@
Let us begin with the estimates for the first term on the right hand side of (5.38). Using (4.10) and the same computation of oscillation error in (4.17), we have for any $t\in[k,k+1]$, $k\in\mathbb{Z}$
$w_{q+1}^{(p)}\otimes w_{q+1}^{(p)}$ (5.39)
@@
Taking the trace on both sides of (5.39) and using the fact that $\mathring{R}_{\ell}$ is traceless, we deduce for $t\in[k,k+1]$, $k\in\mathbb{Z}$
@@
In the following, we estimate each term separately. By the choice of parameters, we find
@@
which requires $b^{2}\beta<1$ and $a$ sufficiently large to absorb the constant. We use (3.8f) on $\mathring{R}_{q}$, $\operatorname{supp}\varphi_{\ell}\subset[0,\ell]$ and $c_{R}<c_{*}$ to obtain for $t\in{\mathbb{R}}$
@@
For the third term in (5.40), using (3.3), (3.7), (3.8b) and (3.8c) we obtain
@@
where we used $14b^{2}\beta+2\beta<$ 1 to have $\lambda_{q}^{-\frac{1}{2}+\beta}<\lambda_{q}^{-2b^{2}\beta}$ and $\lambda_{q}^{-\frac{5}{6}+\beta}<\lambda_{q}^{-2b^{2}\beta}$, and $a$ was chosen sufficiently large to absorb the constant.
Recalling the definition of $\eta_{k,j}$ and $a_{(\xi)}$ in (4.9), we can simplify the last term in (5.40). For any fixed $t$ within the interval $[k,k+1]$, the sum concerning $j$ and $j^{\prime}$ is the finite sum of at most two terms. Hence, applying (5.22) with $n=m$ and using the estimates (5.1), (5.2) and (C.7) we obtain
@@
where we used $b^{3}\gamma+2b^{2}\beta<1$ and $5m+5<bm$ to have $\lambda_{q}^{4m+2+(m+2)b^{3}\gamma}<\lambda_{q}^{bm-2b^{2}\beta}$, and we chose $a$ sufficiently large to absorb the constant. This completes the bound for (5.40).
Going back to (5.38), we have to control
@@
where we used $b\geqslant 7$, $14b^{3}\gamma+4b^{2}\beta<$ 1 to have $\lambda_{q}^{2b^{3}\gamma-\frac{1}{2}}<\lambda_{q}^{-2b^{2}\beta}$, $\lambda_{q}^{4+3b^{3}\gamma}<\lambda_{q}^{b-2b^{2}\beta}$ and $\lambda_{q}^{8+4b^{3}\gamma}<\lambda_{q}^{2b-2b^{2}\beta}$, and $a$ was chosen sufficiently large to absorb the constant. Using (3.3), (3.7), (3.8b), (5.9) and (5.15) we obtain
@@
where we used $2b\beta+\beta<\frac{\gamma\sigma}{8}$ to have $\lambda_{q}^{b\beta+2b^{2}\beta}<\lambda_{q}^{\frac{\gamma\sigma}{8}b}$. We also used $14b^{2}\beta+2b\beta<$ 1, $3b\beta<\gamma$ to have $\lambda_{q}^{b\beta-\frac{1}{2}}<\lambda_{q}^{-2b^{2}\beta}$, $\lambda_{q}^{-b^{2}\gamma+b\beta+b\gamma}<\lambda_{q}^{-2b^{2}\beta}$ and chose $a$ large enough to absorb the constant.
For the last term in (5.41), we recall (cf.(4.15)) that $w_{q+1}$ may be written as the curl of a vector field as
@@
where we used $b\geqslant 7$, $12b^{2}\beta+b^{3}\gamma<$ 1 to have $\lambda_{q}^{b^{3}\gamma+\frac{3}{2}}<\lambda_{q}^{b-2b^{2}\beta}$ and $\lambda_{q}^{b^{3}\gamma+b\gamma}<\lambda_{q}^{b-2b^{2}\beta}$, and $a$ was chosen sufficiently large to absorb the constant. Summing over the above estimates, we obtain
@@
Hence, the proof of Proposition 3.3 is complete. ∎
@@ 6. Stationary solutions to the stochastic Euler equations
We consider that the trajectory space defined in Section 1.1 is $\mathcal{T}=C({\mathbb{R}};C^{\kappa})\times C({\mathbb{R}};C^{\kappa})$ and we take $\kappa=\frac{\vartheta}{2}$ in this section. The corresponding shifts $S_{t}$, $t\in{\mathbb{R}}$, on trajectories are given by
@@
Recall the notion of stationary solution introduced in Definition 1.5. Our main result of this section is the existence of stationary solutions which are constructed as limits of ergodic averages of solutions from Theorem 3.4. This also implies their non-uniqueness.
@@
$\sup_{s\in{\mathbb{R}}}\mathbf{E}\|u(\cdot+s)\|^{2r}_{C^{\beta^{\prime}}([-N,N],C^{\beta^{\prime\prime}}_{x})}\lesssim N.$ (6.2)
Indeed, reviewing the proof of Theorem 3.4 and using interpolation, we deduce for some $\beta^{\prime}\in(0,\frac{\vartheta}{2})$ and $\beta^{\prime\prime}\in(\frac{\vartheta}{2},\vartheta)$ satisfying $\beta^{\prime}+\beta^{\prime\prime}<\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot b^{5}}},\frac{1}{21{\cdot}b^{4}}\}}$
@@
where we used $b\geqslant 7$, $(\frac{1}{4}-\beta^{\prime})(1-\beta^{{}^{\prime\prime}})>\frac{1}{16}$ and $\beta^{\prime\prime}<{\mathord{{\rm min}}}{\{\frac{\sigma}{4b},1\}}$ in the last two inequalities. Hence we obtain $\lim_{q\rightarrow\infty}z_{q}=z$ in $L^{p}(\Omega,C^{\beta^{\prime}}(\mathbb{R};C^{\beta^{\prime\prime}}))$ for any $p\geqslant 1$. Then letting $u=v+z$, we obtain an $(\mathcal{F}_{t})_{t\in\mathbb{R}}$-adapted analytically weak solution $u$ to (1.1) and
@@
Furthermore, by $\|u\|_{C^{\beta^{\prime}}_{[-N,N]}C^{\beta^{\prime\prime}}_{x}}\leqslant\sum_{i=-N}^{N-1}\|u\|_{C^{\beta^{\prime}}_{[i,i+1]}C^{\beta^{\prime\prime}}_{x}}$ for any $N\in\mathbb{N}$ and Minkowski’s inequality, we obtain
$\sup_{s\in{\mathbb{R}}}\left[\mathbf{E}\|u(\cdot+s)\|^{2r}_{C^{\beta^{\prime}}_{[-N,N]}C^{\beta^{\prime\prime}}_{x}}\right]^{\frac{1}{2r}}$ $\leqslant\sup_{s\in{\mathbb{R}}}\sum_{i=-N}^{N-1}\left[\mathbf{E}\|u(\cdot+s)\|^{2r}_{C^{\beta^{\prime}}_{[i,i+1]}C^{\beta^{\prime\prime}}_{x}}\right]^{\frac{1}{2r}}$
$\leqslant 2N\sup_{t\in\mathbb{R}}\left[\mathbf{E}\|u\|^{2r}_{C^{\beta^{\prime}}_{[t,t+1]}C^{\beta^{\prime\prime}}_{x}}\right]^{\frac{1}{2r}}=2N|\!|\!|u|\!|\!|_{C^{\beta^{\prime}}_{t}C^{\beta^{\prime\prime}}_{x},2r}\lesssim N,$
where the implicit constants are independent of $N$. Hence, assertion (6.2) holds.
Now, we define the ergodic averages of the solution $(u,B)$ as the probability measures on the trajectory space ${\mathcal{T}}$
@@
For $R_{N}>0$, $N\in\mathbb{N}$, by Arzelà–Ascoli Theorem and the above claim, we have that the set
@@
is relatively compact in $C({\mathbb{R}};C^{\frac{\vartheta}{2}})$ since $\frac{\vartheta}{2}<\beta^{\prime\prime}$. We then deduce that the time shifts $S_{t}u$, $t\in{\mathbb{R}}$, are tight on $C({\mathbb{R}};C^{\frac{\vartheta}{2}})$. Since $S_{t}B$ is a Wiener process for every $t\in{\mathbb{R}}$, the law of $S_{t}B$ is tight. Accordingly, for any $\varepsilon>0$ there is a compact set $F_{\varepsilon}$ in ${\mathcal{T}}$ such that
$\sup_{t\in{\mathbb{R}}}{\mathbf{P}}(S_{t}(u,B)\in F^{c}_{\varepsilon})<\varepsilon.$
@@
$\nu_{T}(F_{\varepsilon}^{c})=$ $\frac{1}{T}\int_{0}^{T}\mathbf{P}(S_{t}(u,B)\in F^{c}_{\varepsilon}){\mathord{{\rm d}}}t<\varepsilon.$
Therefore, there is a weakly converging subsequence of the probability measures $\nu_{T}$, $T\geqslant 0$. The rest of the proof follows by Jakubowski–Skorokhod representation theorem and the same argument as in [HZZ22b, Theorem 4.1]. ∎
Using the above result and choosing different $K$, we conclude Theorem 1.6.
@@ Appendix 1 A Beltrami waves
In this part, we recall the Beltrami waves from [BV19, Section 5.4] which is adapted to the convex integration scheme in Proposition 3.3. We point out that the construction is entirely deterministic, meaning that none of the functions below depends on $\omega$. Let us begin with the definition of Beltrami waves.
Given $\xi\in\mathbb{S}^{2}\cap\mathbb{Q}^{3}$, let $A_{\xi}\in\mathbb{S}^{2}\cap\mathbb{Q}^{3}$ obey
@@
is $\mathbb{T}^{3}$ periodic, divergence-free, and is an eigenfunction of the $\mathrm{curl}$ operator with eigenvalue $\lambda$. That is, $W_{(\xi)}$ is a complex Beltrami plane wave. The following lemma states a useful property for linear combinations of complex Beltrami plane waves, which can be found in [BV19, Proposition 5.5].
@@
The key point of the construction is that the abundance of Beltrami waves allows us to find several such flows $v$ with the property that
@@
equals a prescribed symmetric matrix $R$. Indeed we will need to select these flows so as to depend smoothly on the matrix $R$, at least when $R$ belongs to a neighborhood of the identity matrix. In view of (A.3), such selection is made possible by the following lemma from [BV19, Proposition 5.6].
@@
$\Lambda_{\alpha}\subset\mathbb{S}^{2}\cap\mathbb{Q}^{3},\qquad\alpha\in\{0,1\},$
@@
$\gamma_{\xi}^{(\alpha)}\in C^{\infty}(B_{c_{*}}(\mathrm{Id})),\qquad\alpha\in\{0,1\},\ \xi\in\Lambda_{\alpha},$